# Federated Causal Inference from Multi-Site Observational Data via Propensity Score Aggregation

### v2

@@ 4 Experiments
We evaluate our proposed Fed-IPW and Fed-AIPW using the MW weights estimated either via federated multinomial logistic regression (well specified in DGP B) or via a two-layer Neural Network (NN). We compare these methods against several competitors: Fed-IPW and Fed-AIPW with the alternative DW weights based on Gaussian density estimation (well specified in DGP A); the Centralized Oracle (Def. 1); meta-analysis IPW/AIPW with sample-size weighting (Meta-SW) (Def. 2); and the one-shot inverse-variance weighted IPW estimator (1S-IVW) of Xiong et al. (2023), evaluated under favorable conditions with shared propensity-score parameters across sites. For all estimators, propensity scores are fit via logistic regression and outcome models via linear regression, implying misspecification in the latter.
Figures 44 (top: DGP A; bottom: DGP B) summarize the results under the three overlap regimes. Before discussing each setting, we highlight several general observations. First, Fed-IPW-DW is unbiased only under DGP A, where the Gaussian specification holds, and exhibits bias under DGP B. By contrast, Fed-IPW-MW adapts across data-generating processes: with logistic membership weights it is unbiased in DGP B, and with a more flexible neural network classifier it is unbiased for both DGP A and DGP A, reflecting the greater modeling flexibility of MW relative to DW’s restrictive density specification. Second, Fed-AIPW enjoys its doubly robust property and remains unbiased across all overlap levels by double robustness; despite misspecified linear outcome models, accurate propensity estimation ensures consistency. Finally, Fed-IPW typically attains lower variance than the centralized oracle IPW, consistent with well-documented efficiency gains from using estimated propensity scores (Hirano et al., 2003).
@@
Figure 2 considers a setting where all local propensity scores share a common subset of 5 of 10 logistic regression coefficients with data generated from DGP B. This setup matches the assumptions of the 1S-IVW method of Xiong et al. (2023), which relies on prior knowledge of the shared parameters to aggregate them—an assumption not required by our method. Fed-MW remains unbiased and attains the lowest variance, matching that of 1S-IVW, even without access to the additional information about shared parameters.
@@
Figure 2: Comparison to Xiong et al. (2023) (IPW).
@@
Real data. We analyze the multi-site Traumabase cohort (Mayer et al., 2020; Colnet et al., 2024) to estimate the effect of tranexamic acid on mortality across $K=14$ centers. The local datasets are highly imbalanced in site sizes (106 to 2,092 patients) and treatment arms (e.g., site 11: 4 treated vs. 121 controls); there are 17 covariates and n=8,248n=8{,}248 patients in total, of whom 638 were treated. Covariates are standardized federatively by sharing site means and variances. We focus on AIPW estimators: local propensities $e_{k}$ are estimated via logistic regression; outcome models $\mu_{1},\mu_{0}$ are trained with FedAvg logistic regression (5,000 rounds, 1 local epoch, step size $\eta=0.1$); MW are learned with a FedAvg neural network (one hidden layer, 128 units), while DW are based on Gaussian density estimation. Competitors include Meta-SW (same local models, computed only on sites with enough treated units—where the number of treated observations exceeds the covariate dimension) and a centralized AIPW benchmark obtained using the R package grf’s probability_forest function (Tibshirani et al., 2018) on the pooled data. Empirical confidence intervals are constructed from 150 bootstrap resamples.
Figure 3 shows that our federated estimators closely match the centralized benchmark and exhibit lower variance than Meta-SW, which in this application departs from the centralized estimate. Among federated methods, Fed-MW is closest to the nonparametric centralized estimator—reflecting its more flexible modeling of propensities and membership—whereas Fed-DW is less stable, likely due to noisy per-site covariance estimates for a 17-dimensional Gaussian ($\approx 17^{2}$ parameters) in small sites.
@@ 5 Conclusion, limitations and future work
We propose a theoretically grounded framework for federated causal inference that leverages membership weights to construct valid pseudo-populations across silos. These weights, estimated via flexible parametric or nonparametric models, improve overlap and yield more stable ATE estimates, without sharing raw data. Our framework accommodates heterogeneous local propensity score estimation strategies, supports external control arms, and remains robust to even extreme local treatment–control imbalances. Although sufficiently large per-site datasets are still needed—particularly in high-dimensional settings—our approach is especially well suited to a moderate number of large silos, where FL most effectively increases effective sample size.
Promising avenues for future work include principled handling of site effects (i.e., relaxing Assumption 2) and extending our framework to CATE estimation. While the ATE remains central in econometrics, biomedicine, and public policy, considering CATE is important for moving towards personalization. Our work lays a foundation for federated CATE estimation: most learners (T-, S-, X-, and R-learners) rely on nuisance quantities such as propensity scores, and our federated estimation framework could be incorporated into these methods, although further work is needed to assess its formal properties and practical performance.
@@ A.1 ATE Identification and Sufficient Conditions
Under Assumptions 1 and 2, the variables in $X_{i}$ are a sufficient adjustment set, so that the ATE is identifiable as
@@
$=\mathbb{E}\left[\mathbb{E}\left[Y(1)\mid X\right]-\mathbb{E}\left[Y(0)\mid X\right]\right]$
$=\mathbb{E}\left[\mathbb{E}\left[Y(1)\mid X,W=1\right]-\mathbb{E}\left[Y(0)\mid X,W=0\right]\right]$ Assumptions 1 and 2
$=\mathbb{E}\left[\mathbb{E}\left[Y\mid X,W=1\right]-\mathbb{E}\left[Y\mid X,W=0\right]\right]$ SUTVA
@@ A.2 Proof of Theorem 1
Unbiasedness. We have:
$\mathbb{E}\left[\hat{\tau}^{\mathrm{IPW^{*}}}\right]$ $=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)\right]$
$=\frac{1}{n}\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{E}\left[\sum_{i=1}^{n}\left(\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\right)\cdot\mathbb{1}_{[H_{i}=k]}\mid n_{1},\dots,n_{K}\right]\right]$
$=\frac{1}{n}\sum_{k=1}^{K}\mathbb{E}\left[n_{k}\right]\left(\mathbb{E}\left[\frac{W_{i}Y_{i}}{e(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\mid H_{i}=k\right]\right)$
$=\sum_{k=1}^{K}\rho_{k}\left(\mathbb{E}\left[\frac{W_{i}Y_{i}}{e(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\mid H_{i}=k\right]\right)$ (6)
Let us focus on the first term of the sum:
$\mathbb{E}\left[\frac{W_{i}Y_{i}}{e(X_{i})}\mid H_{i}=k\right]$ $=\mathbb{E}\left[\frac{W_{i}Y_{i}(1)}{e(X_{i})}\mid H_{i}=k\right]$ SUTVA
$=\mathbb{E}\left[\mathbb{E}\left[\frac{W_{i}Y_{i}(1)}{e(X_{i})}\mid X_{i},H_{i}=k\right]\right]$
$=\mathbb{E}\left[\frac{\mathbb{E}\left[W_{i}\mid X_{i},H_{i}=k\right]\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]}{e(X_{i})}\right]$ Assumptions 1 and 2
$=\mathbb{E}\left[\frac{e_{k}(X_{i})\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]}{e(X_{i})}\right]$
$=\mathbb{E}\left[\frac{e_{k}(X_{i})\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]}{\mathbb{E}\left[e_{k}(X_{i})\mid X_{i},H_{i}=k\right]}\right]$ Definition of $e(X_{i})$
$=\mathbb{E}\left[\mathbb{E}\left[Y_{i}(1)\mid X_{i},H_{i}=k\right]\mathbb{E}\left[\frac{e_{k}(X_{i})}{\mathbb{E}\left[e_{k}(X_{i})\right]}\mid X_{i}\right]\right]$
$=\mathbb{E}\left[Y_{i}(1)\mid H_{i}=k\right]$
Similarly, we have $\mathbb{E}\left[\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}\mid H_{i}=k\right]=\mathbb{E}\left[Y_{i}(0)\mid H_{i}=k\right]$, so that $\mathbb{E}\left[\hat{\tau}^{\mathrm{IPW^{*}}}\right]=\sum_{k=1}^{K}\rho_{k}\tau_{k}$, which concludes the proof of unbiasedness of the oracle multi-site centralized IPW estimator.
For the AIPW estimator, following the same steps as in section˜A.2, we have
$\mathbb{E}\left[\hat{\tau}^{\mathrm{AIPW^{*}}}\right]$ $=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\left(\mu_{1}(X_{i})-\mu_{0}(X_{i})+\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\right)\right]$
$=\sum_{k=1}^{K}\rho_{k}\Bigg(\mathbb{E}\left[\mu_{1}(X_{i})-\mu_{0}(X_{i})\mid H_{i}=k\right]$
$\quad\quad+\mathbb{E}\left[\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}\mid H_{i}=k\right]-\mathbb{E}\left[\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}\mid H_{i}=k\right]\Bigg)$
$=\sum_{k=1}^{K}\rho_{k}\left(\mathbb{E}\left[\tau(X_{i})\mid H_{i}=k\right]+0+0\right)$
$=\tau$
@@ B.2 Federated Training Procedure
1:Input: $K$ sites, $E$ local steps, $\eta$ learning rate, $T$ communication rounds, $B$ batch size
2:Initialize global parameters $\Theta_{0}$
3:for $t=1$ to $T$ do
4:  for each client $k\in[1,\dots,K]$ in parallel do
5:   $\Theta_{t+1}^{(k)}\leftarrow$ LocalUpdate($k,\Theta_{t}$)
6:  end for
7:  $\Theta_{t+1}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{n}\Theta_{t+1}^{(k)}$ // FedAvg aggregation
8:end for
9:LocalUpdate($k$, $\Theta$):
10:for $e=1$ to $E$ do
11:  Sample mini-batch $\mathcal{B}_{k}$ of size $B$ from local data $\mathcal{D}_{k}$
12:  Compute predictions $\omega_{k}(X_{i};\Theta)$ for $i\in\mathcal{B}_{k}$
13:  Compute gradient $\nabla\ell(\Theta;\mathcal{B}_{k})$ of cross-entropy loss
14:  $\Theta\leftarrow\Theta-\eta\nabla\ell(\Theta;\mathcal{B}_{k})$
15:end for
16:return $\Theta$
@@ Appendix C Simulation Details
The parameters common to all settings are shown in Table 1, where $\gamma_{2}^{\mathrm{(weak)}}=[-2.5,-1,-0.15,-0.15,0,-0.15,-1,-0.15,-0.15,0]$ and $\gamma_{2}^{\mathrm{(good)}}=[-.05,-.1,-.05,-.1,.05,-.1,-.05,-.1,.05,-.1]$.
@@
$\gamma_{k}$ $\begin{array}[]{c}[-.25,.25,\\ -.25,-.25,.25,\\ -.25,-.25,.25,\\ -.25,.25]\end{array}$
No overlap: not logistic
(only controls)
Weak overlap: $\gamma_{2}^{\mathrm{(weak)}}$
Good overlap: $\gamma_{2}^{\mathrm{(good)}}$
$\begin{array}[]{c}[.15,-.15,\\ .15,-.15,.15,\\ -.15,.15,-.15,\\ .15,-.15]\end{array}$

### v3

@@ 4.1 Synthetic Data
We evaluate our proposed Fed-IPW and Fed-AIPW using the MW weights estimated either via federated multinomial logistic regression (well specified in DGP B) or via a two-layer Neural Network (NN). We compare these methods against several competitors: Fed-IPW and Fed-AIPW with the alternative DW weights based on Gaussian density estimation (well specified in DGP A); the Centralized Oracle (Def. 1); meta-analysis IPW/AIPW with sample-size weighting (Meta-SW) (Def. 3); and the one-shot inverse-variance weighted IPW estimator (1S-IVW) of Xiong et al. (2023), evaluated under favorable conditions with shared propensity-score parameters across sites. For all estimators, propensity scores are fit via logistic regression and outcome models via linear regression, implying misspecification in the latter.
Behavior under varying overlap. Figures 44 (top: DGP A; bottom: DGP B) summarize the results under the three overlap regimes. Before discussing each setting, we highlight several general observations. First, Fed-IPW-DW is unbiased only under DGP A, where the Gaussian specification holds, and exhibits bias under DGP B. By contrast, Fed-IPW-MW adapts across data-generating processes: with logistic membership weights it is unbiased in DGP B, and with a more flexible neural network classifier it is unbiased for both DGP A and DGP A, reflecting the greater modeling flexibility of MW relative to DW’s restrictive density specification. Second, Fed-AIPW enjoys its doubly robust property and remains unbiased across all overlap levels by double robustness; despite misspecified linear outcome models, accurate propensity estimation ensures consistency. Finally, Fed-IPW typically attains lower variance than the centralized oracle IPW, consistent with well-documented efficiency gains from using estimated propensity scores (Hirano et al., 2003).
@@
Comparison to Xiong et al. (2023). Figure 3 considers a setting where all local propensity scores share a common subset of 5 of 10 logistic regression coefficients with data generated from DGP B. This setup matches the assumptions of the 1S-IVW method of Xiong et al. (2023), which relies on prior knowledge of the shared parameters to aggregate them—an assumption not required by our method. Fed-MW remains unbiased and attains the lowest variance, matching that of 1S-IVW, even without access to the additional information about shared parameters.
Handling site effets. We assess the benefit of adjusting for site effects when site-level covariates $Z$ are observed (Section˜3.3). We simulate $K=30$ sites, with heterogeneity arising both from diverse distributions of individual-level covariates $X$ (via covariate shift induced by $H\sim\mathrm{Logit}(\theta^{\top}X)$) and from site-level variation in outcomes captured by $Z$. Outcomes follow $Y=X^{\top}\beta+Z^{\top}\delta+\tau W+\varepsilon$, so that $H$ affects $Y$ only through $Z$. Federated methods incorporate $Z$ only in the membership weights, while meta-analysis adjusts only for $X$ as it naturally handles site effects. As shown in Table 1, incorporating $Z$ into our method eliminates confounding due to both covariate shift ($H\to X$) and mediated site effects ($H\to Z\to Y$). As a result, our Federated IPW/AIPW markedly improves MSE, especially for small $n_{k}$, where meta-analysis suffers from unstable local nuisance estimation.
@@
Estimator $n_{k}{=}20$ $n_{k}{=}60$ $n_{k}{=}100$
IPW Oracle $(e^{\star}(X,Z))$ 0.853 0.443 0.091
IPW Meta-SW $(\hat{e}(X))$ 5.176 0.777 0.590
IPW Fed-MW (Logit, $\hat{e}(X,Z)$) 0.427 0.132 0.043
IPW Fed-MW (Logit, $\hat{e}(X)$) 2.001 1.384 0.327
IPW Fed-MW (NN, $\hat{e}(X,Z)$) 1.083 0.465 0.206
IPW Fed-MW (NN, $\hat{e}(X)$) 2.100 1.418 0.356
AIPW Oracle $(e^{\ast}(X,Z))$ 0.015 0.005 0.002
IPW Meta-SW $(\hat{e}(X))$ 10.673 0.346 0.084
AIPW Fed-MW (Logit, $\hat{e}(X,Z)$) 0.022 0.004 0.002
AIPW Fed-MW (Logit, $\hat{e}(X)$) 2.026 1.348 0.323
AIPW Fed-MW (NN, $\hat{e}(X,Z)$) 0.020 0.004 0.002
AIPW Fed-MW (NN, $\hat{e}(X)$) 2.043 1.349 0.315
Table 1: Mean Squared Error (MSE) of federated and meta-analysis estimators under Assumption 5.
Figure 2: Comparison to Xiong et al. (2023) (IPW estimators).
@@
4.2 Real Data
We analyze the multi-site Traumabase cohort (Mayer et al., 2020; Colnet et al., 2024) to estimate the effect of tranexamic acid on mortality across K=14K\!=\!14 centers. The local datasets are highly imbalanced in sizes (106 to 2,092 patients) and treatment arms (e.g., site 11: 4 treated vs. 121 controls); there are 17 covariates and n=8,248n=8{,}248 patients in total, of whom 638 were treated. Covariates are standardized federatively by sharing site means and variances. We focus on AIPW estimators: local propensities $e_{k}$ are estimated using the R package grf’s probability_forest function (Tibshirani et al., 2018) on large sites, and logistic regressions on smaller ones; outcome models $\mu_{1},\mu_{0}$ are trained with FedAvg logistic regression (5,000 rounds, 1 local epoch, step size $\eta=0.1$); MW are learned with a FedAvg neural network (NN, one hidden layer, 128 units) and a federated logistic regression. We also report a centralized random forest (RF) MW as a strong benchmark. DW are based on Gaussian density estimation. Competitors include Meta-SW (same local models, computed only on sites with enough treated units—where the number of treated observations exceeds the covariate dimension) and a centralized AIPW benchmark obtained using probability_forest function on the pooled data. Empirical confidence intervals are constructed from 150 bootstrap resamples.
Figure 3 shows that our federated MW estimators closely match the centralized benchmark and exhibit lower variance than Meta-SW, which in this application departs from the centralized estimate. Among federated methods, Fed-MW are closest to the nonparametric centralized estimators—reflecting their more flexible modeling of propensities and membership—whereas Fed-DW is less stable, likely due to noisy per-site covariance estimates for a 17-dimensional Gaussian ($\approx 17^{2}$ parameters) in small sites.
@@ 5 Conclusion, Limitations and Future work
We propose a theoretically grounded framework for federated causal inference that leverages membership weights to construct valid causal estimates across silos. These weights, estimated via flexible parametric or nonparametric models, improve overlap and yield more stable ATE estimates. Our framework accomodates arbitrary heterogeneity in local propensity score functions, does not require local overlap to hold, enabling the inclusion of small sites and external control arms, without sharing individual-level data. Our approach can also be naturally extended to handle some form of site effects. Although sufficiently large per-site datasets are still needed—particularly in high-dimensional settings—to learn local propensity scores accurately, our approach is especially well suited to a moderate number of large silos, where federated learning most effectively increases effective sample size.
A promising avenue for future work is to extend our framework to CATE estimation. While the ATE remains central in econometrics, biomedicine, and public policy, considering CATE is important for moving towards personalization. Our work lays a foundation for federated CATE estimation: most learners (T-, S-, X-, and R-learners) rely on nuisance quantities such as propensity scores, and our federated estimation framework could be incorporated into these methods, although further work is needed to assess its formal properties and practical performance.
@@ A.1 ATE Identification and Sufficient Conditions
Under Assumptions 1 and 2, the variables in $X$ form a sufficient adjustment set. We identify the ATE as follows:
@@
$=\mathbb{E}\left[\mathbb{E}\left[Y(1)\mid X,H\right]-\mathbb{E}\left[Y(0)\mid X,H\right]\right]$ Law of Iterated Expectations
$=\mathbb{E}\left[\mathbb{E}\left[Y(1)\mid X,H,W=1\right]-\mathbb{E}\left[Y(0)\mid X,H,W=0\right]\right]$ Assumption 1
$=\mathbb{E}\left[\mathbb{E}\left[Y(1)\mid X,W=1\right]-\mathbb{E}\left[Y(0)\mid X,W=0\right]\right]$ Assumption 2 and Lemma 4.3 Dawid [1979]
$=\mathbb{E}\left[\mathbb{E}\left[Y\mid X,W=1\right]-\mathbb{E}\left[Y\mid X,W=0\right]\right]$ Consistency
@@ A.2 Proof of Theorem 1
Unbiasedness. We show the unbiasedness of the first term of the IPW estimator. The derivation for the second term is symmetric.
$\mathbb{E}\left[\frac{WY}{e(X)}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[\frac{WY(1)}{e(X)}\middle|X,H\right]\right]$ SUTVA
$=\mathbb{E}\left[\frac{\mathbb{E}\left[W\mid X,H\right]\mathbb{E}\left[Y(1)\mid X,H\right]}{e(X)}\right]$ Assumption 1
$=\mathbb{E}\left[\frac{e_{H}(X)\mathbb{E}\left[Y(1)\mid X,H\right]}{e(X)}\right]$ $\text{Definition of local score }e_{H}(X)$
$=\mathbb{E}\left[\frac{e_{H}(X)\mathbb{E}\left[Y(1)\mid X\right]}{e(X)}\right]$ Assumption 2
$=\mathbb{E}_{X}\left[\frac{\mathbb{E}\left[Y(1)\mid X\right]}{e(X)}\mathbb{E}_{H}\left[e_{H}(X)\mid X\right]\right]$ $\text{Iterated Expectation over }H$
$=\mathbb{E}_{X}\left[\frac{\mathbb{E}\left[Y(1)\mid X\right]}{e(X)}\sum_{k=1}^{K}\mathbb{P}(H=k\mid X)e_{k}(X)\right]$
$=\mathbb{E}_{X}\left[\frac{\mathbb{E}\left[Y(1)\mid X\right]}{e(X)}e(X)\right]$ $\text{Definition of global }e(X)$
$=\mathbb{E}\left[\mathbb{E}[Y(1)\mid X]\right]=\mathbb{E}[Y(1)]$
Note that the step applying Assumption 2 relies on the consequence that $\mathbb{E}[Y(1)\mid X,H]=\mathbb{E}[Y(1)\mid X]$ (which is derived from Assumption 2 combined with Assumption 1).
Similarly, we have $\mathbb{E}\left[\frac{(1-W)Y}{1-e(X)}\right]=\mathbb{E}[Y(0)]$, so that $\mathbb{E}\left[\hat{\tau}^{\mathrm{IPW^{*}}}\right]=\tau$, which concludes the proof of unbiasedness of the oracle multi-site centralized IPW estimator.
For the AIPW estimator, we note that $\mathbb{E}[W(Y-\mu_{1}(X))/e(X)]=0$ following the same logic as above (replacing $Y(1)$ with the residual $Y(1)-\mu_{1}(X)$ which has mean zero given $X$). Thus:
$\mathbb{E}\left[\hat{\tau}^{\mathrm{AIPW^{*}}}\right]$ $=\mathbb{E}\left[\mu_{1}(X)-\mu_{0}(X)+\frac{W(Y-\mu_{1}(X))}{e(X)}-\frac{(1-W)(Y-\mu_{0}(X))}{1-e(X)}\right]$
$=\mathbb{E}[\mu_{1}(X)-\mu_{0}(X)]+0-0$
$=\mathbb{E}[\tau(X)]=\tau$
@@ B.2 Federated Training Procedure
0:$K$ sites, $E$ local steps, $\eta$ learning rate, $T$ communication rounds, $B$ batch size
1: Initialize global parameters $\Theta_{0}$
2:for $t=1$ to $T$ do
3:  for each client $k\in[1,\dots,K]$ in parallel do
4:   $\Theta_{t+1}^{(k)}\leftarrow$ LocalUpdate($k,\Theta_{t}$)
5:  end for
6:  $\Theta_{t+1}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{n}\Theta_{t+1}^{(k)}$ // FedAvg aggregation
7:end for
8:Function LocalUpdate($k$, $\Theta$):
9:for $e=1$ to $E$ do
10:  Sample mini-batch $\mathcal{B}_{k}$ of size $B$ from local data $\mathcal{D}_{k}$
11:  Compute predictions $\omega_{k}(X_{i};\Theta)$ for $i\in\mathcal{B}_{k}$
12:  Compute gradient $\nabla\ell(\Theta;\mathcal{B}_{k})$ of cross-entropy loss
13:  $\Theta\leftarrow\Theta-\eta\nabla\ell(\Theta;\mathcal{B}_{k})$
14:end for
15:Return $\Theta$
@@
Note that this federated procedure remains valid for learning $H|X$ despite each site $k$ locally holding only one label ($H=k$).
@@ Appendix C Simulation Details
The parameters common to all settings are shown in Table 2, where
$\gamma_{2}^{\mathrm{(weak)}}=[-2.5,-1,-0.15,-0.15,0,-0.15,-1,-0.15,-0.15,0]$
and
$\gamma_{2}^{\mathrm{(good)}}=[-.05,-.1,-.05,-.1,.05,-.1,-.05,-.1,.05,-.1]$
@@
$\gamma_{k}$ $\begin{array}[]{c}[-.25,.25,\\ -.25,-.25,.25,\\ -.25,-.25,.25,\\ -.25,.25]\end{array}$ No overlap: not logistic (only controls) Weak overlap: $\gamma_{2}^{\mathrm{(weak)}}$ Good overlap: $\gamma_{2}^{\mathrm{(good)}}$ $\begin{array}[]{c}[.15,-.15,\\ .15,-.15,.15,\\ -.15,.15,-.15,\\ .15,-.15]\end{array}$
@@
Real data. We classified a site as ‘large’ if it had at least $5\times d$ observations in each treatment arm (minimum 85 per arm in our setting). For such sites, we estimated $e_{k}$ with probability forests with $2,000$ trees; otherwise, we used logistic regression. We estimated Fed-MW RF with a centralized random forest for $P(H|X)$ with $500$ trees.